# Rule-Based Gillespie Simulation of Chemical Systems

### v1

@@ 6.1 Example 1: Lotka-Volterra System
The reaction rate callback will assign a constant rate of $100$ to replication_rabbitsreplication\\ \_rabbits, and a rate of $0.1$ to replication_foxesreplication\_foxes. For the growth and death reactions, we employ the input and output rate callbacks: a constant input flow of 55 only for $Y$; and an output flow of 55 for both species $Y$ and $Z$. This way we open the system. Regarding initial concentrations we define $[Y]=100$ and $[Z]=10$. We let the simulation run until reaching $150,000$ iterations. As shown in Figure 2, the resulting concentration evolution shows expected oscillatory behavior, characteristic of the Lotka-Volterra system. This confirms the model’s capacity to represent predator-prey dynamics accurately within an open system.
@@ 6.3 Example 3: Conditional Callbacks
Given a simple assembly system of molecules denoted AA and $B$:
@@
We set the initial concentrations to $[A]=100$ and $[B]=500$ molecules. Regarding the reaction rate callback, the bindingbinding rule will have a rate constant of $0.01$. For the unbindingunbinding rule, the rate constant will be 11 if $[A]>200$; and $0.01$ otherwise—balancing the rates of creation and destruction of bonds between AA and $B$. Additionally, the system will have an input flow of AA that we set to 55 while $[A]<200$, producing a rapid increase; and to $0.01$ otherwise.
We can visualize the resulting evolution of concentrations in Figure 6. We observe an initial decay in the concentration of $B$, as this molecule is in fixed amounts and it’s being used to produce ABAB. Similarly, AA is consumed very quickly at the beginning. When almost all $B$ molecules are consumed, the rate of production of ABAB stabilizes close to $500$ copies, around $t=80$. At this point, the bindingbinding and unbindingunbinding rules are less reactive, and we can observe a rapid increase in the concentration of AA. By design, when $A]\geq 200$, at approximately $t=125$, the input flow of AA is reduced significantly, and the unbindingunbinding reaction will kinetically dominate over the bindingbinding one, resulting in equilibrium for all the species.

### v2

@@ 6.1 Example 1: Lotka-Volterra System
The reaction rate callback will assign a constant rate of $100$ to replication_rabbitsreplication\\ \_rabbits, and a rate of $0.1$ to replication_foxesreplication\_foxes. For the growth and death reactions, we employ the input and output rate callbacks: a constant input flow of $5$ only for $Y$; and an output flow of $5$ for both species $Y$ and $Z$. This way we open the system. Regarding initial concentrations we define $[Y]=100$ and $[Z]=10$. We let the simulation run until reaching $150,000$ iterations. As shown in Figure 2, the resulting concentration evolution shows expected oscillatory behavior, characteristic of the Lotka-Volterra system. This confirms the model’s capacity to represent predator-prey dynamics accurately within an open system.
@@ 6.3 Example 3: Conditional Callbacks
Given a simple assembly system of molecules denoted $A$ and $B$:
@@
We set the initial concentrations to $[A]=100$ and $[B]=500$ molecules. Regarding the reaction rate callback, the bindingbinding rule will have a rate constant of $0.01$. For the unbindingunbinding rule, the rate constant will be $1$ if $[A]>200$; and $0.01$ otherwise—balancing the rates of creation and destruction of bonds between $A$ and $B$. Additionally, the system will have an input flow of $A$ that we set to $5$ while $[A]<200$, producing a rapid increase; and to $0.01$ otherwise.
We can visualize the resulting evolution of concentrations in Figure 6. We observe an initial decay in the concentration of $B$, as this molecule is in fixed amounts and it’s being used to produce $AB$. Similarly, $A$ is consumed very quickly at the beginning. When almost all $B$ molecules are consumed, the rate of production of $AB$ stabilizes close to $500$ copies, around $t=80$. At this point, the bindingbinding and unbindingunbinding rules are less reactive, and we can observe a rapid increase in the concentration of $A$. By design, when $A]\geq 200$, at approximately $t=125$, the input flow of $A$ is reduced significantly, and the unbindingunbinding reaction will kinetically dominate over the bindingbinding one, resulting in equilibrium for all the species.